# A general approach to asymptotic elimination of aggregation functions and generalized quantifiers

### v3

@@ 1. Introduction
The use of (generalized) quantifiers is a way to increase the expressivity of a logic, beyond the use of functions (often called connectives) which, for some $k$, assign a truth value to any $k$-tuple of truth values. In the context of artificial intelligence and machine learning it makes sense to use aggregation functions, such as for example the average of a sequence of numbers, to increase expressivity, for example when defining probability distributions, or when defining queries. (See e.g. [3, pages 10, 117, 382] or [5, pages 31, 54].) In the context of this article, an aggregation function is a function that takes, for some integer $k>0$, a $k$-tuple $(\bar{p}_{1},\ldots,\bar{p}_{k})$ as input, where each $\bar{p}_{i}$ is a finite sequence of reals from the unit interval, and gives a real number in the unit interval as output. In addition, for each $i$, the order of the reals in $\bar{p}_{i}$ must not matter for the output value. (Therefore some people prefer to view each $\bar{p}_{i}$ as a finite multiset.) Note that we allow $\bar{p}_{i}$ and $\bar{p}_{j}$ to have different length (if $i\neq j$). To systematize the use of aggregation functions one can incorporate them into the syntax and semantics of a formal logic, just as done with quantifiers. But even if a sequence of numbers contains only the numbers ‘0’ and ‘1’ the average of the sequence may well be a number strictly between 0 and 1. So if we use aggregation functions in a formal logic we have to accept that the (truth) values of formulas can be other numbers than 0 or 1.
@@
In this work we consider a quite general (formal) logic, which we call $PLA^{*}$ (“probability logic with aggregation functions”), with truth values in the unit interval $[0,1]$ and which uses all possible aggregation functions. (The ‘***’ in $PLA^{*}$ is there to indicate that it is another, more general and flexible, version of the logics $PLA$ and $PLA^{+}$ considered in [14] and [15].) The restriction of truth values to $[0,1]$ is partly motivated by the fact that it is often natural to see such truth values as probabilities, and partly because a sequence of nonnegative numbers can be “normalized” so that the resulting sequence contains only numbers in the unit interval but preserves the variation in the original sequence. Example 2.9 illustrates the expressivity of $PLA^{*}$ by showing that every “stage” (in the iterative approximation) of the PageRank [2] is definable in $PLA^{*}$. The first-order quantifiers $\exists$ and $\forall$ are not lost by using aggregation functions instead of quantifiers, because $\exists$ and $\forall$ can be expressed by using the aggregation functions ‘maximum’ and ‘minimum’ (see Remark 2.10). Moreover, as we will see in Proposition 3.3, every condition which can be expressed by first-order logic extended by any generalized quantifiers in the sense of Mostowski can be expressed by a formula in $PLA^{*}$.
@@
We now consider the following context: $\sigma$ is a finite and relational signature (vocabulary), $D_{n}$, $n=1,2,3,\ldots$, is a sequence of finite domains (i.e. sets) such that, for every $n$, the cardinality of $D_{n+1}$ is greater than the cardinality of $D_{n}$, $\mathbf{W}_{n}$ is a set of $\sigma$-structures with domain $D_{n}$, and $\mathbb{P}_{n}$ is a probability distribution on $\mathbf{W}_{n}$. Results have been proven, for various 2-valued logics, various sets of structures $\mathbf{W}_{n}$, and various sequences of probability distributions $\mathbb{P}_{n}$, which say that all (or some) formulas of the logic are, with high probability, equivalent to a formula without quantifiers (or that this is not the case). Examples are [4, 6, 7, 9, 11, 12, 13]. Often such a result can be used to prove a so-called logical convergence law, or even a zero-one law. This kind of result often has implications for the expressivity of the logic considered and for the existence of computationally efficient ways of estimating, on a large domain, the probability of a query defined by the logic.
@@
Within the specified context we can ask, for $\varphi(\bar{x})\in PLA^{*}$, $0\leq c<d\leq 1$, and $\bar{a}\in(D_{n})^{|\bar{x}|}$, what the probability is that the value of $\varphi(\bar{a})$ in a random $\mathcal{A}\in\mathbf{W}_{n}$ is in the interval $(c,d)$. Let us call a formula of any logic aggregation-free if it contains no aggregation function and no quantifier. For the reasons mentioned above, it is of interest to understand under what conditions for a $PLA^{*}$-formula $\varphi(\bar{x})$ there is a $PLA^{*}$-formula $\psi(\bar{x})$ such that $\psi(\bar{x})$ is aggregation-free, or at least of lesser complexity than $\varphi$, and for all $\varepsilon>0$ the probability that the values of the two formulas differ by at most $\varepsilon$ (for all substitutions of parameters for the free variables $\bar{x}$) tends to 1 as $n\to\infty$; in this case we say that $\varphi(\bar{x})$ and $\psi(\bar{x})$ are asymptotically equivalent (see Definition 5.2).
@@
In general, the first step seems to be intricately connected to the kind of structures and the kind of probability distributions considered. But the second step can be formulated in such a way that it makes sense in a wide variety of situations, as we will show in Assumption 5.5 and Theorem 5.9.. We hope that Theorem 5.9 can be used several times in the future once its preconditions (Assumption 5.5) have been shown. This avoids carrying out a similar proof again and again, as already done in [14] and [15], and also in e.g. [12] and [13], although it is less clear in the later cases.
Theorem 5.9 also give us a general understanding of conditions under which aggregation functions can be asymptotically eliminated from $PLA^{*}$-formulas. Indirectly it also gives us an understanding of when Mostowski style generalized quantifiers can be asymptoticaly eliminated. Hence our Theorem 5.9 is similar in spirit, but different in the technical approach, to some results of Kaila [11, Theorems 4.4 and 4.7].
@@
Section 4 focuses on two continuity properties of aggregation functions that are crucial to this study. The first continuity property has he advantage that it is relatively easy to check whether an aggregation function has it. The other property is formulated so that it is easy to use in the proof of Theorem 5.9. However, as stated by Proposition 4.10, both “local” continuity properties are very closely related, possibly equivalent although we have not been able to show this. The “global” versions of the two continuity properties are indeed equivalent.
@@ 2. Logic with aggregation functions
In this section we define the logic that we will work with, called $PLA^{*}$. It is an extension of the logics $PLA$ and $PLA^{+}$ considered in [14] and [15], respectively, by allowing a more flexible use of aggregation functions. Formulas of $PLA^{*}$ may take any (truth) value in the unit interval $[0,1]$ and we think of the values 0 and 1 as corresponding to “false” and “true”. $PLA^{*}$ uses aggregation functions instead of quantifiers to “aggregate” over a domain. In Section 3 we will see that any condition expressed by a formula of first-order logic extended by generalized quantifiers in the sense of Mostowski [18] can be expressed by a formula of $PLA^{*}$. Recall that $[0,1]^{<\omega}$ denotes the set of all finite nonempty sequences of reals in the unit interval $[0,1]$.
@@
    For each $c\in[0,1]$, $c\in PLA^{*}(\sigma)$ (i.e. $c$ is a formula) and $Fv(c)=\emptyset$. We also let $\bot$ and $\top$ denote 00 and $1$, respectively.
@@
(i) When denoting a formula in $PLA^{*}(\sigma)$ by for example $\varphi(\bar{x})$ then we assume that $\bar{x}$ is a sequence of different variables and that every variable in the formula denoted by $\varphi(\bar{x})$ belongs to $\mathrm{rng}(\bar{x})$ (but we do not require that every variable in $\mathrm{rng}(\bar{x})$ actually occurs in the formula).
@@
(Semantics of $PLA^{*}(\sigma)$) For every $\varphi\in PLA^{*}(\sigma)$ and every sequence of distinct variables $\bar{x}$ such that $Fv(\varphi)\subseteq\mathrm{rng}(\bar{x})$ we associate a mapping from pairs $(\mathcal{A},\bar{a})$, where $\mathcal{A}$ is a finite $\sigma$-structure and $\bar{a}\in A^{|\bar{x}|}$, to $[0,1]$. The number in $[0,1]$ to which $(\mathcal{A},\bar{a})$ is mapped is denoted $\mathcal{A}(\varphi(\bar{a}))$ and is defined by induction on the complexity of formulas.
@@
(ii) If $\varphi(\bar{x})\in PLA^{*}(\sigma)$, $\mathcal{A}$ is a finite $\sigma$-structure and that $\bar{a}\in A^{|\bar{x}|}$, then ‘$\mathcal{A}\models\varphi(\bar{a})$’ means the same as ‘$\mathcal{A}\big{(}\varphi(\bar{a})\big{)}=1$’.
@@
$PR_{0}(x)=1/N\text{ where $N$ is the number of sites,}$,
@@
Suppose that $PR_{k}(x)$ is expressed by $\varphi_{k}(x)\in PLA^{*}(\sigma)$. Note that multiplication is a continuous connective from $[0,1]^{2}$ to $[0,1]$ so it can be used in $PLA^{*}(\sigma)$-formulas. Then observe that the quantity $|OUT_{y}|^{-1}$ is expressed by the $PLA^{*}(\sigma)$-formula
@@
(Using $\exists$ and $\forall$ as abbreviations) Motivated by the discussion in Remark 2.10, if $\varphi(\bar{x},\bar{y})\in PLA^{*}(\sigma)$ is a formula which can only take the values 0 or 1 (e.g a boolean combination of atomic first-order formulas) then we may use ‘$\exists\bar{y}\varphi(\bar{x},\bar{y})$’ to mean the same as ‘$\max(\varphi(\bar{x},\bar{y}):\bar{y}:\top)$’, and ‘$\forall\bar{y}\varphi(\bar{x},\bar{y})$’ to mean the same as ‘$\min(\varphi(\bar{x},\bar{y}):\bar{y}:\top)$’.
@@ 3. Expressing generalized quantifiers with aggregation functions
Proof. For any $\varphi(\bar{x})\in FOGQ(\sigma)$, finite $\sigma$-structure $\mathcal{A}$ and $\bar{a}\in A^{|\bar{x}|}$, let ‘$\mathcal{A}\big{(}\varphi(\bar{a})\big{)}=1$’ mean the same as ‘$\mathcal{A}\models\varphi(\bar{a})$’ and let ’$\mathcal{A}\big{(}\varphi(\bar{a})\big{)}=0$’ mean the same as ‘$\mathcal{A}\not\models\varphi(\bar{a})$’.
In Remark 2.10 we saw that the proposition holds if $\varphi(\bar{x})$ is a first-order formula. In the same remark we also saw that $\neg$, $\wedge$, $\vee$ and $\to$ can, with Lukasiewicz semantics, be expressed by continuous functions from $[0,1]$ or $[0,1]^{2}$ to $[0,1]$, and $\exists$ and $\forall$ can be expressed by the aggregation functions max and min, respectively.
@@
$\begin{cases}1\qquad\text{ if }(D,X_{1},\ldots,X_{k})\in Q\text{ where $D=[m]$, $m=\max(m_{1},\ldots,m_{k})$}\\ \qquad\quad\text{ and }X_{i}=\{j\in[m_{i}]:r_{i,j}=1\},\text{ and}\\ 0\qquad\text{ otherwise.}\end{cases}$
@@ 4. Notions of continuity for aggregation functions
Then we show (see Proposition 4.10) that ct-continuity and up-continuity are closely related. Nevertheless we consider both notions useful because ct-continuity (with respect to some parameters) is usually easier to verify for concrete aggregation functions, while up-continuity is tailor made for making the proof of “elimination of an aggregation function” (Theorem 5.9) work out.
@@ 4.2. Uniform point continuity
It follows that $\mu_{1}^{u}$ and $\mu_{\infty}^{o}$ are symmetric and satisfy the triangle inequality so they are pseudometrics on $[0,1]^{<\omega}$. None of them is a metric since it can happen that $\mu_{1}^{u}(\bar{p},\bar{q})=0$ and $\bar{p}\neq\bar{q}$. For example, if $\bar{p}=(0,1/2,1)$ and $\bar{q}=(0,0,1/2,1/2,1,1)$ then $\mu_{1}^{u}(\bar{p},\bar{q})=0$. Also observe that for all $\bar{p},\bar{q}\in[0,1]^{<\omega}$, $\mu_{1}^{u}(\bar{p},\bar{q}),\mu_{\infty}^{o}(\bar{p},\bar{q})\leq 1$.
@@
    $\text{which equal $c_{i,j}$}\big{\}}.$
@@
We only show that the first limit is 0, since the the second limit is treated in the same way. We will see that this is a consequence of Condition (2) of Definition 4.8 of up-continuity. By the choice of $\bar{p}_{n}$ (as convergence testing) and construction of $\bar{p}^{\prime}_{n}$ we have $|\bar{p}_{n}|=|\bar{p}^{\prime}_{n}|$, $\lim_{n\to\infty}|\bar{p}_{n}|=\infty$, $\mathrm{rng}(\bar{p}^{\prime}_{n})\subseteq\{c_{1},\ldots,c_{k}\}$, and for every $\delta>0$ there is $n_{0}$ such that, $\mu_{\infty}^{o}(\bar{p}_{n},\bar{p}^{\prime}_{n})<\delta$ if $n>n_{0}$ (because $I_{1},\ldots,I_{k}$ can be chosen with diameter at most $\delta$). From (4.1) and the construction of $\bar{p}^{\prime}_{n}$ it follows that for every $\delta>0$ there is $n_{0}$ such that for all $n>n_{0}$ and $j=1,\ldots,k$, the number of coordinates in $\bar{p}^{\prime}_{n}$ which equal $c_{j}$ is between $(\alpha_{j}-\delta)|\bar{p}^{\prime}_{n}|$ and $(\alpha_{j}+\delta)|\bar{p}^{\prime}_{n}|$. As $F$ is up-continuous with respect to $c_{1},\ldots,c_{k}$ and $\alpha_{1},\ldots,\alpha_{k}$, it follows from condition (2) of Definition 4.8 that $\lim_{n\to\infty}|F(\bar{p}_{n})-F(\bar{p}^{\prime}_{n})|=0$.
@@
$\text{which equal $c_{j}$}\big{\}}.$
@@
Define $\bar{p}^{\prime}_{1}=\bar{p}_{1,1}$, $\bar{q}^{\prime}_{1}=\bar{q}_{1,1}$, and for $n>1$, $\bar{p}^{\prime}_{n}=\bar{p}_{n,N}$ and $\bar{q}^{\prime}_{n}=\bar{q}_{n,N}$ where $N$ is larger than $|\bar{p}^{\prime}_{m}|$ and $|\bar{q}^{\prime}_{m}|$ for all $m<n$. Then $|\bar{p}^{\prime}_{n}|<|\bar{p}^{\prime}_{n+1}|$, $|\bar{q}^{\prime}_{n}|<|\bar{q}^{\prime}_{n+1}|$, $\mu_{1}^{u}(\bar{p}^{\prime}_{n},\bar{q}^{\prime}_{n})<1/n$ and $|F(\bar{p}^{\prime}_{n})-F(\bar{q}^{\prime}_{n})|>\varepsilon$ for all $n$. Let $\bar{p}^{\prime}_{n,i}$ denote the $i$th coordinate of $\bar{p}^{\prime}$. Since $[0,1]^{k}$ is a compact topological space, the sequence of $k$-tuples
(4.2) (|$\big{(}|\{i:\bar{p}^{\prime}_{n,i}=c_{1}\}|/|\bar{p}^{\prime}_{n}|,\ldots,|\{i:\bar{p}^{\prime}_{n,i}=c_{k}\}|/|\bar{p}^{\prime}_{n}|\big{)}$
@@
From Proposition 4.4 and Proposition 4.10 it follows that the aggregation functions ‘am’ (arithmetic mean) and ‘gm’ (geometric mean) are up-continuous with respect to all parameters, and that the aggregation functions ‘max’ and ‘min’ are, for all $m,k_{1},\ldots,k_{m}\in\mathbb{N}^{+}$, all $c_{i,j}\in[0,1]$ and all $\alpha_{i,j}\in(0,1]$, up-continuous with respect to the parameters $c_{i,j}$ and $\alpha_{i,j}$. From earlier examples and Proposition 4.10 it follows that ‘$\mathrm{length}^{-\beta}$’ (where $\beta\in(0,1)$) and ‘tsum’ (“truncated sum”) are up-continuous with respect to all parameters. More examples of aggregation functions that are ct-continuous and up-continuous with respect to all parameters, or possibly with the requirement that $\alpha_{i,j}>0$ for all $i$ and $j$, are found in [14] and [15].
In the next two examples we use the following notation: If $\alpha\in[0,1]$ and $\bar{p}\in[0,1]^{<\omega}$, then #$\#(\alpha,\bar{p})$ is the number of coordinates in $\bar{p}$ which are equal to $\alpha$. The remaining two examples consider aggregation functions obtained from generalized quantifiers, as discussed in Section 3, and show that rather mild conditions on a sequence of parameters imply that such aggregation functions are ct-continuous, respectively up-continuous, with respect to that sequence of parameters.
@@
$F(\bar{p})=\begin{cases}1\qquad\text{ if }\#(1,\bar{p})/|\bar{p}|\geq\beta,\\ 0\qquad\text{ otherwise.}\end{cases}$

### v4

@@ 1. Introduction
The use of (generalized) quantifiers is a way to increase the expressivity of a logic, beyond the use of functions (often called connectives) which, for some $k$, assign a truth value to any $k$-tuple of truth values. In the context of artificial intelligence and machine learning it makes sense to use aggregation functions, such as for example the average of a sequence of numbers, to increase expressivity, for example when defining probability distributions, or when defining queries. (See e.g. [3, pages 10, 117, 382] or [5, pages 31, 54].) In the context of this article, an aggregation function is a function that takes, for some integer $k>0$, a $k$-tuple $(\bar{p}_{1},\ldots,\bar{p}_{k})$ as input, where each $\bar{p}_{i}$ is a finite sequence of reals from the unit interval, and gives a real number in the unit interval as output. In addition, for each $i$, the order of the reals in $\bar{p}_{i}$ should not matter for the output value. (Therefore some people prefer to view each $\bar{p}_{i}$ as a finite multiset.) Note that we allow $\bar{p}_{i}$ and $\bar{p}_{j}$ to have different length (if $i\neq j$). To systematize the use of aggregation functions one can incorporate them into the syntax and semantics of a formal logic, just as done with quantifiers. But even if a sequence of numbers contains only the numbers ‘0’ and ‘1’ the average of the sequence may well be a number strictly between 0 and 1. So if we use aggregation functions in a formal logic we have to accept that the (truth) values of formulas can be other numbers than 0 or 1.
@@
In this work we consider a quite general (formal) logic, which we call $PLA^{*}$ (“probability logic with aggregation functions”), with truth values in the unit interval $[0,1]$ and which uses all possible aggregation functions. (The ‘*’ in $PLA^{*}$ is there to indicate that it is another, more general and flexible, version of the logics $PLA$ and $PLA^{+}$ considered in [17] and [18].) The restriction of truth values to $[0,1]$ is partly motivated by the fact that it is often natural to see such truth values as degrees of uncertainty (e.g. probabilities), and partly because a sequence of nonnegative numbers can be “normalized” so that the resulting sequence contains only numbers in the unit interval but preserves the variation in the original sequence. Example 2.9 illustrates the expressivity of $PLA^{*}$ by showing that every “stage” (in the iterative approximation) of the PageRank [2] is definable in $PLA^{*}$. The first-order quantifiers $\exists$ and $\forall$ are not lost by using aggregation functions instead of quantifiers, because $\exists$ and $\forall$ can be expressed by using the aggregation functions ‘maximum’ and ‘minimum’ (see Remark 2.10). Moreover, as we will see in Proposition 3.3, every condition which can be expressed by first-order logic extended by any generalized quantifiers in the sense of Mostowski [21, 22] can be expressed by a formula in $PLA^{*}$. (Other early work on generalized quantifiers include Lindström [19] and Hajek, Havel, and Chytil [9] about generating plausible hypotheses in the early days of artificial intelligence.)
@@
We now consider the following context: $\sigma$ is a finite and relational signature (vocabulary), $D_{n}$, $n=1,2,3,\ldots$, is a sequence of finite domains (i.e. sets) such that, for every $n$, the cardinality of $D_{n+1}$ is greater than the cardinality of $D_{n}$, $\mathbf{W}_{n}$ is a set of $\sigma$-structures with domain $D_{n}$, and $\mathbb{P}_{n}$ is a probability distribution on $\mathbf{W}_{n}$. Results have been proven, for various 2-valued logics, various sets of structures $\mathbf{W}_{n}$, and various sequences of probability distributions $\mathbb{P}_{n}$, which say that all (or some) formulas of the logic are, with high probability, equivalent to a formula without quantifiers (or that this is not the case). Examples include [4, 6, 7, 10, 12, 13, 14]. Often such a result can be used to prove a so-called logical convergence law, or even a zero-one law. This kind of result often has implications for the expressivity of the logic considered and for the existence of computationally efficient ways of estimating, on a large domain, the probability of a query defined by the logic.
@@
Within the specified context we can ask, for $\varphi(\bar{x})\in PLA^{*}$, $0\leq c<d\leq 1$, and $\bar{a}\in(D_{n})^{|\bar{x}|}$, what the probability is that the value of $\varphi(\bar{a})$ in a random $\mathcal{A}\in\mathbf{W}_{n}$ is in the interval $(c,d)$. Let us call a formula of any logic aggregation-free if it contains no aggregation function and no quantifier. For the reasons mentioned above, it is of interest to understand under what conditions for a $PLA^{*}$-formula $\varphi(\bar{x})$ there is a $PLA^{*}$-formula $\psi(\bar{x})$ such that $\psi(\bar{x})$ is aggregation-free, or at least of smaller complexity than $\varphi$, and for all $\varepsilon>0$ the probability that the values of the two formulas differ by at most $\varepsilon$ (for all substitutions of parameters for the free variables $\bar{x}$) tends to 1 as $n\to\infty$; in this case we say that $\varphi(\bar{x})$ and $\psi(\bar{x})$ are asymptotically equivalent (see Definition 5.2).
@@
In general, the first step seems to be intricately connected to the kind of structures and the kind of probability distributions considered. But the second step can be formulated in such a way that it makes sense in a wide variety of situations, as we will show in Assumption 5.5 and Theorem 5.9. We hope that Theorem 5.9 can be used several times in the future once its preconditions (Assumption 5.5) have been shown to hold. This avoids carrying out a similar proof again and again, as already done in [17] and [18], and also in e.g. [13] and [14], although it is less clear in the latter cases. In fact, Theorem 5.9 has been used in precisely this way in [15] and [16] (which were written after the submission of this article).
Theorem 5.9 also gives us a general understanding of conditions under which aggregation functions can be asymptotically eliminated from $PLA^{*}$-formulas. Indirectly it also gives us an understanding of when Mostowski style generalized quantifiers can be asymptoticaly eliminated. Hence our Theorem 5.9 is similar in spirit, but different in the technical approach, to some results of Kaila [12, Theorems 4.4 and 4.7].
@@
Section 4 focuses on two continuity properties of aggregation functions that are crucial to this study. The first continuity property has the advantage that it is relatively easy to check whether an aggregation function has it. The other property is formulated so that it is easy to use in the proof of Theorem 5.9. However, as stated by Proposition 4.10, both “local” continuity properties are very closely related, possibly equivalent although we have not been able to show this. The “global” versions of the two continuity properties are indeed equivalent.
@@ 2. Logic with aggregation functions
In this section we define the logic that we will work with, called $PLA^{*}$. It is an extension of the logics $PLA$ and $PLA^{+}$ considered in [17] and [18], respectively, by allowing a more flexible use of aggregation functions. Formulas of $PLA^{*}$ may take any (truth) value in the unit interval $[0,1]$ and we think of the values 0 and 1 as corresponding to “false” and “true”. $PLA^{*}$ uses aggregation functions instead of quantifiers to “aggregate” over a domain. In Section 3 we will see that any condition expressed by a formula of first-order logic extended by generalized quantifiers in the sense of Mostowski [22] can be expressed by a formula of $PLA^{*}$. Recall that $0,1]^{<\omega}$ denotes the set of all finite nonempty sequences of reals in the unit interval $[0,1]$.
@@
    For each $c\in[0,1]$, $c\in PLA^{*}(\sigma)$ (i.e. $c$ is a formula) and $Fv(c)=\emptyset$. We also let $\bot$ and $\top$ denote 0 and $1$, respectively.
@@
(i) When denoting a formula in $PLA^{*}(\sigma)$ by e.g. $\varphi(\bar{x})$ we assume that $\bar{x}$ is a sequence of different variables and that every variable in the formula denoted by $\varphi(\bar{x})$ belongs to $\mathrm{rng}(\bar{x})$ (but we do not require that every variable in $\mathrm{rng}(\bar{x})$ actually occurs in the formula).
@@
(Semantics of $PLA^{*}(\sigma)$) For every $\varphi\in PLA^{*}(\sigma)$ and every sequence of distinct variables $\bar{x}$ such that $Fv(\varphi)\subseteq\mathrm{rng}(\bar{x})$ we associate a mapping from pairs $(\mathcal{A},\bar{a})$, where $\mathcal{A}$ is a finite $\sigma$-structure and $\bar{a}\in A^{|\bar{x}|}$, to $[0,1]$. The number in $[0,1]$ to which $(\mathcal{A},\bar{a})$ is mapped is denoted by $\mathcal{A}(\varphi(\bar{a}))$ and is defined by induction on the complexity of formulas.
@@
(ii) If $\varphi(\bar{x})\in PLA^{*}(\sigma)$, $\mathcal{A}$ is a finite $\sigma$-structure and that $\bar{a}\in A^{|\bar{x}|}$, then $\mathcal{A}\models\varphi(\bar{a})$’ means the same as $\mathcal{A}\big(\varphi(\bar{a})\big)=1$’.
@@
$PR_{0}(x)=1/N\text{ where $N$ is the number of sites,}$
@@
Suppose that $PR_{k}(x)$ is expressed by $\varphi_{k}(x)\in PLA^{*}(\sigma)$. Note that multiplication is a continuous connective from $0,1]^{2}$ to $[0,1]$ so it can be used in $PLA^{*}(\sigma)$-formulas. Then observe that the quantity $|OUT_{y}|^{-1}$ is expressed by the $PLA^{*}(\sigma)$-formula
@@
(Using $\exists$ and $\forall$ as abbreviations) Motivated by the discussion in Remark 2.10, if $\varphi(\bar{x},\bar{y})\in PLA^{*}(\sigma)$ is a formula which can only take the values 0 or 1 (e.g a boolean combination of atomic first-order formulas) then we may use $\exists\bar{y}\varphi(\bar{x},\bar{y})$’ to mean the same as $\max(\varphi(\bar{x},\bar{y}):\bar{y}:\top)$’, and $\forall\bar{y}\varphi(\bar{x},\bar{y})$’ to mean the same as $\min(\varphi(\bar{x},\bar{y}):\bar{y}:\top)$’.
@@ 3. Expressing generalized quantifiers with aggregation functions
Proof. For any $\varphi(\bar{x})\in FOGQ(\sigma)$, finite $\sigma$-structure $\mathcal{A}$ and $\bar{a}\in A^{|\bar{x}|}$, let $\mathcal{A}\big(\varphi(\bar{a})\big)=1$’ mean the same as $\mathcal{A}\models\varphi(\bar{a})$’ and let $\mathcal{A}\big(\varphi(\bar{a})\big)=0$’ mean the same as $\mathcal{A}\not\models\varphi(\bar{a})$’.
In Remark 2.10 we saw that the proposition holds if $\varphi(\bar{x})$ is a first-order formula. In the same remark we also saw that $\neg$, $\wedge$, $\vee$ and $\to$ can, with Lukasiewicz semantics, be expressed by continuous functions from $[0,1]$ or $0,1]^{2}$ to $[0,1]$, and $\exists$ and $\forall$ can be expressed by the aggregation functions max and min, respectively.
@@
[…]
@@ 4. Notions of continuity for aggregation functions
Then we show (see Proposition 4.10) that ct-continuity and up-continuity are closely related. Nevertheless we consider both notions useful because ct-continuity (with respect to some parameters) is usually easier to verify for concrete aggregation functions, while up-continuity is tailored for making the proof of “elimination of an aggregation function” (Theorem 5.9) work out.
@@ 4.2. Uniform point continuity
It follows that $\mu_{1}^{u}$ and $\mu_{\infty}^{o}$ are symmetric and satisfy the triangle inequality so they are pseudometrics on $0,1]^{<\omega}$. None of them is a metric since it can happen that $\mu_{1}^{u}(\bar{p},\bar{q})=0$ and $\bar{p}\neq\bar{q}$. For example, if $\bar{p}=(0,1/2,1)$ and $\bar{q}=(0,0,1/2,1/2,1,1)$ then $\mu_{1}^{u}(\bar{p},\bar{q})=0$. Also observe that for all $\bar{p},\bar{q}\in[0,1]^{<\omega}$, $\mu_{1}^{u}(\bar{p},\bar{q}),\mu_{\infty}^{o}(\bar{p},\bar{q})\leq 1$.
@@
    $\text{which equals $c_{i,j}$}\big\}.$
@@
We only show that the first limit is 0, since the the second limit is treated in the same way. We will see that this is a consequence of Condition (2) of Definition 4.8 of up-continuity. By the choice of $\bar{p}_{n}$ (as convergence testing) and construction of $\bar{p}^{\prime}_{n}$ we have $|\bar{p}_{n}|=|\bar{p}^{\prime}_{n}|$, $\lim_{n\to\infty}|\bar{p}_{n}|=\infty$, $\mathrm{rng}(\bar{p}^{\prime}_{n})\subseteq\{c_{1},\ldots,c_{k}\}$, and for every $\delta>0$ there is $n_{0}$ such that, $\mu_{\infty}^{o}(\bar{p}_{n},\bar{p}^{\prime}_{n})<\delta$ if $n>n_{0}$ (because $I_{1},\ldots,I_{k}$ can be chosen with diameter at most $\delta$). From (4.1) and the construction of $\bar{p}^{\prime}_{n}$ it follows that for every $\delta>0$ there is $n_{0}$ such that for all $n>n_{0}$ and $j=1,\ldots,k$, the number of coordinates in $\bar{p}^{\prime}_{n}$ which equals $c_{j}$ is between $(\alpha_{j}-\delta)|\bar{p}^{\prime}_{n}|$ and $(\alpha_{j}+\delta)|\bar{p}^{\prime}_{n}|$. As $F$ is up-continuous with respect to $c_{1},\ldots,c_{k}$ and $\alpha_{1},\ldots,\alpha_{k}$, it follows from condition (2) of Definition 4.8 that $\lim_{n\to\infty}|F(\bar{p}_{n})-F(\bar{p}^{\prime}_{n})|=0$.
@@
$\text{which equals $c_{j}$}\big\}.$
@@
Define $\bar{p}^{\prime}_{1}=\bar{p}_{1,1}$, $\bar{q}^{\prime}_{1}=\bar{q}_{1,1}$, and for $n>1$, $\bar{p}^{\prime}_{n}=\bar{p}_{n,N}$ and $\bar{q}^{\prime}_{n}=\bar{q}_{n,N}$ where $N$ is larger than $|\bar{p}^{\prime}_{m}|$ and $|\bar{q}^{\prime}_{m}|$ for all $m<n$. Then $|\bar{p}^{\prime}_{n}|<|\bar{p}^{\prime}_{n+1}|$, $|\bar{q}^{\prime}_{n}|<|\bar{q}^{\prime}_{n+1}|$, $\mu_{1}^{u}(\bar{p}^{\prime}_{n},\bar{q}^{\prime}_{n})<1/n$ and $|F(\bar{p}^{\prime}_{n})-F(\bar{q}^{\prime}_{n})|>\varepsilon$ for all $n$. Let $\bar{p}^{\prime}_{n,i}$ denote the $i$th coordinate of $\bar{p}^{\prime}$. Since $0,1]^{k}$ is a compact topological space, the sequence of $k$-tuples
(4.2) $\big(|\{i:\bar{p}^{\prime}_{n,i}=c_{1}\}|/|\bar{p}^{\prime}_{n}|,\ldots,|\{i:\bar{p}^{\prime}_{n,i}=c_{k}\}|/|\bar{p}^{\prime}_{n}|\big)$
@@
From Proposition 4.4 and Proposition 4.10 it follows that the aggregation functions ‘am’ (arithmetic mean) and ‘gm’ (geometric mean) are up-continuous with respect to all parameters, and that the aggregation functions ‘max’ and ‘min’ are, for all $m,k_{1},\ldots,k_{m}\in\mathbb{N}^{+}$, all $c_{i,j}\in[0,1]$ and all $\alpha_{i,j}\in(0,1]$, up-continuous with respect to the parameters $c_{i,j}$ and $\alpha_{i,j}$. From earlier examples and Proposition 4.10 it follows that $\mathrm{length}^{-\beta}$’ (where $\beta\in(0,1)$) and ‘tsum’ (“truncated sum”) are up-continuous with respect to all parameters. More examples of aggregation functions that are ct-continuous and up-continuous with respect to all parameters, or possibly with the requirement that $\alpha_{i,j}>0$ for all $i$ and $j$, are found in [17] and [18].
In the next two examples we use the following notation: If $\alpha\in[0,1]$ and $\bar{p}\in[0,1]^{<\omega}$, then $(\alpha,\bar{p})$ is the number of coordinates in $\bar{p}$ which are equal to $\alpha$. The remaining two examples consider aggregation functions obtained from generalized quantifiers, as discussed in Section 3, and show that rather mild conditions on a sequence of parameters imply that such aggregation functions are ct-continuous, respectively up-continuous, with respect to that sequence of parameters.
@@
[…]